An ice skating arena charges an admission fee for each child plus a rental fee for each pair of ice skates. John paid the admission fees for his six nephews and rented five pairs of ice skates. He was charged $32.00. Juanita paid the admission fees for her seven grandchildren and rented five pairs of ice skates. She was charged $35.25. What is the admission fee? What is the rental fee for a pair of skates? Write a system of equations to find the admission and rental fees. Identify the variables, write the equations, and solve the system using Elimination or Substitution methods. Write complete Answers. Show all your work step by step. Write a short explanation about why you chose the best method to solve the problem.
Answer by checkley79(3341)

6F+5S=32
7F+5S=35.25 SUBTRACT.
--------------------------
-F=-3.25
F=$3.25 ANS. FOR THE FEE PER CHILD.
6*3.25+5S=32
19.50+5S=32
5S=32-19.50
5S=12.50
S=12.50/5
S=$2.50 ANS. FOR THE COST OF SKATE RENTAL.
PROOF:
7*3.25+5*2.50=35.25
22.75+12.50=35.25
35.25=35.25